# Viologen-based supramolecular crystal gels: gelation kinetics and sensitivity to temperature

### v1

@@ IV Conclusion
In conclusion, this study sheds light on the formation and characteristics of viologen-based supramolecular crystal gels, a subset of molecular gels formed by self-assembly of low molecular weight gelators into interconnected crystalline fibers, resulting in a three-dimensional soft solid network. Our in-depth analyzes of the gelation kinetics triggered by thermal quenching reveal the tunability of the 3D network and provide clear insights into the mechanism of gel formation.
Experimental analyses, encompassing UV-Vis spectroscopy, rheology, microscopy, and scattering experiments, provide a comprehensive understanding of the system’s kinetics. We first confirm that the viologen-based gelator $Chol_{2}V^{2+}$ crystallizes into nanometer-radius hollow tubes, which further assemble into micro to millimetric spherulites. Then, we demonstrate the spherulite growth is consistent with the Avrami theory. Local characterization of the growth kinetics using optical microscopy and USALS techniques allowed us to determine independently the exponents in the Avrami model (Eq. 2: $n=n_{N}+dn_{G}$). It reveals that the crystallization process is driven by pre-existing nuclei ($n_{N}=0$), with growth controlled by interfaces ($n_{G}=1$) and characterized by branching resulting in the formation of fractal spherulites of fractal dimension $d$. Finally, our results demonstrate the possibility of adjusting the gel properties by varying the quenching temperature. Lowering the quenching temperature leads to the formation of denser and smaller spherulites. We establish that the elastic properties of the gel cannot be described by the model proposed in [2, 3, 64]. We speculate that the elasticity of the gel stems from both the spherulite elasticity and the way they connect to one another: the spherulites densification obtained at low quenching temperature occurs at the expense of the connectivity between spherulite. This comprehensive analysis provides valuable insights into the design and manipulation of supramolecular crystal gels.
@@ V.2 Radius of the spherulites over time
In order to characterize the formation of ${Chol_{2}V^{2+}/I^{-}}$ CT complexes in the absence of any other anionic species the viologen salt $VI_{2}$ was prepared using the following protocol. Compound $VBr_{2}$ (0.410 g, 0.32 mmol), $LiTFSI$ (3.494 g, 6.4 mmol) and 400 mL technical $EtOH$ were introduced in a 1L round bottom flask and refluxed overnight under stirring.  80% of the solvent was evaporated under reduced pressure, 200mL water were added and the mix was left in the fridge overnight. The mix was filtered on and washed with 3x40mL water. The white solid was then dried under vacuum, 420mg of colorless solid was recovered. This solid was dissolved in 20 mL acetone, Tetrabuthylammonium iodide (820 mg, 2.2mmol) were added and the mix was stirred at room temperature for 1h. The mix was filtered and washed with acetone (3x20mL). $Chol_{2}VI_{2}$ was obtained as an orange solid in 73% yield. The evolution of the UV-Vis absorption spectrum of $Chol_{2}VI_{2}$ is displayed in Fig. 9(a). In Fig. 9(b), we observe that the absorption at $\lambda=440$ nm increase linearly with $Chol_{2}VI_{2}$ concentration leading to an adsorption coefficient is $\epsilon_{\lambda=450nm}^{path=1cm}=6.3\pm 0.3\times 10^{3}$ mol/L/cm.
Figure 9: a) Absorption spectra obtained at 20°C from $VI_{2}$ solutions and gels in pentanol (2mm quartz cuvette). Samples were prepared after applying a heating cooling cycle to $VI_{2}$ suspensions in 1-Pentanol. b) Evolution of the Absorbance at 440 nm with the concentration in $VI_{2}$. The continuous red line to the best linear fit.
@@ Acknowledgment
  • Weiss and Terech [2006] R. G. Weiss and P. Terech, Molecular gels, Materials with Self-Assembled Fibrillar Networks  (2006).
  • Liu and Sawant [2001] X. Y. Liu and P. D. Sawant, Formation kinetics of fractal nanofiber networks in organogels, Applied Physics Letters 79, 3518 (2001).
  • Liu and Sawant [2002] X. Y. Liu and P. D. Sawant, Determination of the fractal characteristic of nanofiber-network formation in supramolecular materials, ChemPhysChem 3, 374 (2002).
  • Lescanne et al. [2003] M. Lescanne, A. Colin, O. Mondain-Monval, F. Fages, and J.-L. Pozzo, Structural aspects of the gelation process observed with low molecular mass organogelators, Langmuir 19, 2013 (2003).
  • Li and Liu [2010] J. L. Li and X. Y. Liu, Architecture of supramolecular soft functional materials: From understanding to micro-/nanoscale engineering, Advanced Functional Materials 20, 3196 (2010).
  • Andrieux et al. [2023] V. Andrieux, T. Gibaud, J. Bauland, T. Divoux, S. Manneville, S. Guy, A. Bensalah-Ledoux, L. Guy, F. Chevallier, D. Frath, and C. Bucher, Chiral and conductive viologen-based supramolecular gels exhibiting tunable charge-transfer properties, J. Mater. Chem. C 11, 12764 (2023).
  • Pernetti et al. [2007] M. Pernetti, K. van Malssen, D. Kalnin, and E. Flöter, Structuring edible oil with lecithin and sorbitan tri-stearate, Food Hydrocolloids 21, 855 (2007).
  • Webber and Langer [2017] M. J. Webber and R. Langer, Drug delivery by supramolecular design, Chemical Society Reviews 46, 6600 (2017).
  • Saunders and Ma [2019] L. Saunders and P. X. Ma, Self-healing supramolecular hydrogels for tissue engineering applications, Macromolecular bioscience 19, 1800313 (2019).
@@
  • Ajayaghosh et al. [2008] A. Ajayaghosh, V. K. Praveen, and C. Vijayakumar, Organogels as scaffolds for excitation energy transfer and light harvesting, Chemical Society Reviews 37, 109 (2008).
  • Kartha et al. [2012] K. K. Kartha, S. S. Babu, S. Srinivasan, and A. Ajayaghosh, Attogram sensing of trinitrotoluene with a self-assembled molecular gelator, Journal of the american chemical society 134, 4834 (2012).
  • Van Bommel et al. [2003] K. J. Van Bommel, A. Friggeri, and S. Shinkai, Organic templates for the generation of inorganic materials, Angewandte Chemie International Edition 42, 980 (2003).
  • Carretti et al. [2010] E. Carretti, M. Bonini, L. Dei, B. H. Berrie, L. V. Angelova, P. Baglioni, and R. G. Weiss, New frontiers in materials science for art conservation: responsive gels and beyond, Accounts of chemical research 43, 751 (2010).
  • Jones and Steed [2016] C. D. Jones and J. W. Steed, Gels with sense: Supramolecular materials that respond to heat, light and sound (2016).
  • Striepe and Baumgartner [2017] L. Striepe and T. Baumgartner, Viologens and their application as functional materials, Chemistry–A European Journal 23, 16924 (2017).
  • Kathiresan et al. [2021] M. Kathiresan, B. Ambrose, N. Angulakshmi, D. E. Mathew, D. Sujatha, and A. M. Stephan, Viologens: a versatile organic molecule for energy storage applications, Journal of Materials Chemistry A 9, 27215 (2021).
  • Fang et al. [2018] W. Fang, Y. Zhang, J. Wu, C. Liu, H. Zhu, and T. Tu, Recent advances in supramolecular gels and catalysis, Chemistry – An Asian Journal 13, 712 (2018).
  • Dzhardimalieva et al. [2020] G. I. Dzhardimalieva, B. C. Yadav, S. Singh, and I. E. Uflyand, Self-healing and shape memory metallopolymers: state-of-the-art and future perspectives, Dalton Trans. 49, 3042 (2020).
  • Chen et al. [2021] H. Chen, V. Brasiliense, J. Mo, L. Zhang, Y. Jiao, Z. Chen, L. O. Jones, G. He, Q.-H. Guo, X.-Y. Chen, et al., Single-molecule charge transport through positively charged electrostatic anchors, Journal of the American Chemical Society 143, 2886 (2021).
  • Li et al. [2021] J. Li, S. Pudar, H. Yu, S. Li, J. S. Moore, J. Rodríguez-López, N. E. Jackson, and C. M. Schroeder, Reversible switching of molecular conductance in viologens is controlled by the electrochemical environment, The Journal of Physical Chemistry C 125, 21862 (2021).
  • Nguyen et al. [2018] Q. V. Nguyen, P. Martin, D. Frath, M. L. Della Rocca, F. Lafolet, S. Bellinck, P. Lafarge, and J.-C. Lacroix, Highly efficient long-range electron transport in a viologen-based molecular junction, Journal of the American Chemical Society 140, 10131 (2018).
  • Han et al. [2020] Y. Han, C. Nickle, Z. Zhang, H. P. Astier, T. J. Duffin, D. Qi, Z. Wang, E. Del Barco, D. Thompson, and C. A. Nijhuis, Electric-field-driven dual-functional molecular switches in tunnel junctions, Nature materials 19, 843 (2020).
  • [23] K. V. Rao, K. Jayaramulu, T. K. Maji, and S. J. George, Supramolecular hydrogels and high-aspect-ratio nanofibers through charge-transfer-induced alternate coassembly, Angewandte Chemie International Edition 49, 4218.
  • Datta et al. [2017] S. Datta, N. Dey, and S. Bhattacharya, Electrochemical probing of hydrogelation induced by the self-assembly of a donor–acceptor complex comprising pyranine and viologen, Chem. Commun. 53, 2371 (2017).
  • Yuan et al. [2017] T. Yuan, Y. Xu, C. Zhu, Z. Jiang, H.-J. Sue, L. Fang, and M. A. Olson, Tunable thermochromism of multifunctional charge-transfer-based supramolecular materials assembled in water, Chemistry of Materials 29, 9937 (2017).
  • [26] T. Yuan, M. Vazquez, A. N. Goldner, Y. Xu, R. Contrucci, M. A. Firestone, M. A. Olson, and L. Fang, Versatile thermochromic supramolecular materials based on competing charge transfer interactions, Advanced Functional Materials 26, 8604.
  • Dhiman et al. [2020] S. Dhiman, K. Jalani, and S. J. George, Redox-mediated, transient supramolecular charge-transfer gel and ink, ACS Applied Materials & Interfaces 12, 5259 (2020).
  • Xue et al. [2004] P. Xue, R. Lu, D. Li, M. Jin, C. Bao, Y. Zhao, and Z. Wang, Rearrangement of the aggregation of the gelator during sol-gel transcription of a dimeric cholesterol-based viologen derivative into fibrous silica, Chem. Mater. 16, 3702 (2004).
  • Suzuki et al. [2001] M. Suzuki, C. C. Waraksa, H. Nakayama, K. Hanabusa, M. Kimura, and H. Shirai, Supramolecular assemblies formed by new l-lysine derivatives of viologens, Chem. Commun. , 2012 (2001).
  • Datta and Bhattacharya [2016] S. Datta and S. Bhattacharya, Carbon-nanotube-mediated electrochemical transition in a redox-active supramolecular hydrogel derived from viologen and an l-alanine-based amphiphile, Chemistry – A European Journal 22, 7524 (2016).
  • Kahlfuss et al. [2018] C. Kahlfuss, T. Gibaud, S. Denis-Quanquin, S. Chowdhury, G. Royal, F. Chevallier, E. Saint-Aman, and C. Bucher, Redox-Induced Molecular Metamorphism Promoting a Sol/Gel Phase Transition in a Viologen-Based Coordination Polymer, Chemistry - A European Journal 24, 13009 (2018).
  • Chowdhury et al. [2021] S. Chowdhury, Q. Reynard-Feytis, C. Roizard, D. Frath, F. Chevallier, C. Bucher, and T. Gibaud, Light-Controlled Aggregation and Gelation of Viologen-Based Coordination Polymers, Journal of Physical Chemistry B 125, 12063 (2021)arXiv:2106.15688 .
  • Roizard et al. [2022] C. Roizard, V. Andrieux, S. A. Shehimy, S. Chowdhury, Q. Reynard-Feytis, C. Kahlfuss, E. Saint-Aman, F. Chevallier, C. Bucher, T. Gibaud, and D. Frath, Photoredox processes in the aggregation and gelation of electron-responsive supramolecular polymers based on viologen, ECS Advances 1, 020502 (2022).
  • Yu et al. [2015] R. Yu, N. Lin, W. Yu, and X. Y. Liu, Crystal networks in supramolecular gels: Formation kinetics and mesoscopic engineering principles, CrystEngComm 17, 7986 (2015).
@@
  • Avrami [1941] M. Avrami, Granulation, phase change, and microstructure kinetics of phase change. iii, The Journal of chemical physics 9, 177 (1941).
  • Shirzad and Viney [2023] K. Shirzad and C. Viney, A critical review on applications of the avrami equation beyond materials science, Journal of the Royal Society Interface 20, 20230242 (2023).
  • Gila-Vilchez et al. [2021] C. Gila-Vilchez, M. C. Mañas-Torres, J. A. González-Vera, F. Franco-Montalban, J. A. Tamayo, F. Conejero-Lara, J. M. Cuerva, M. T. Lopez-Lopez, A. Orte, and L. Á. de Cienfuegos, Insights into the co-assemblies formed by different aromatic short-peptide amphiphiles, Polymer Chemistry 12, 6832 (2021).
  • Ma et al. [2008] Z. Ma, G. Zhang, X. Zhai, L. Jin, X. Tang, M. Yang, P. Zheng, and W. Wang, Fractal crystal growth of poly (ethylene oxide) crystals from its amorphous monolayers, Polymer 49, 1629 (2008).
  • Taguchi et al. [2001] K. Taguchi, H. Miyaji, K. Izumi, A. Hoshino, Y. Miyamoto, and R. Kokawa, Growth shape of isotactic polystyrene crystals in thin films, Polymer 42, 7443 (2001).
  • Heijna et al. [2007] M. C. Heijna, M. J. Theelen, W. J. van Enckevort, and E. Vlieg, Spherulitic growth of hen egg-white lysozyme crystals, The Journal of Physical Chemistry B 111, 1567 (2007).
  • Mareau and Prud’Homme [2005] V. H. Mareau and R. E. Prud’Homme, In-situ hot stage atomic force microscopy study of poly ($\varepsilon$-caprolactone) crystal growth in ultrathin films, Macromolecules 38, 398 (2005).
  • Zhu et al. [2007] D.-S. Zhu, Y.-X. Liu, E.-Q. Chen, M. Li, C. Chen, Y.-H. Sun, A.-C. Shi, R. M. Van Horn, and S. Z. Cheng, Crystal growth mechanism changes in pseudo-dewetted poly (ethylene oxide) thin layers, Macromolecules 40, 1570 (2007).
  • Shi et al. [2009] J. H. Shi, X. Y. Liu, J. L. Li, C. S. Strom, and H. Y. Xu, Spherulitic networks: from structure to rheological property, The journal of physical chemistry b 113, 4549 (2009).
  • Wang et al. [2006] R. Wang, X.-Y. Liu, J. Xiong, and J. Li, Real-time observation of fiber network formation in molecular organogel: supersaturation-dependent microstructure and its related rheological property, The journal of physical chemistry b 110, 7275 (2006).
  • Shtukenberg et al. [2012] A. G. Shtukenberg, Y. O. Punin, E. Gunn, and B. Kahr, Spherulites, Chemical reviews 112, 1805 (2012).
  • Liu et al. [2002] X. Y. Liu, P. D. Sawant, W. B. Tan, I. Noor, C. Pramesti, and B. Chen, Creating new supramolecular materials by architecture of three-dimensional nanocrystal fiber networks, Journal of the American Chemical Society 124, 15055 (2002).
  • Lakshminarayanan et al. [2021] V. Lakshminarayanan, C. Chockalingam, E. Mendes, and J. H. van Esch, Gelation kinetics-structure analysis of ph-triggered low molecular weight hydrogelators, ChemPhysChem 22, 2256 (2021).
  • Lam et al. [2010] R. Lam, L. Quaroni, T. Pedersen, and M. A. Rogers, A molecular insight into the nature of crystallographic mismatches in self-assembled fibrillar networks under non-isothermal crystallization conditions, Soft Matter 6, 404 (2010).
  • Nasr et al. [2021] P. Nasr, H. Leung, F.-I. Auzanneau, and M. A. Rogers, Supramolecular fractal growth of self-assembled fibrillar networks, Gels 7, 46 (2021).
  • Newbloom et al. [2012] G. M. Newbloom, K. M. Weigandt, and D. C. Pozzo, Electrical, mechanical, and structural characterization of self-assembly in poly(3-hexylthiophene) organogel networks, Macromolecules 45, 3452 (2012).
  • Thureau et al. [2021] A. Thureau, P. Roblin, and J. Pérez, Biosaxs on the swing beamline at synchrotron soleil, Journal of Applied Crystallography 54, 1698 (2021).
  • Piau et al. [1999] J.-M. Piau, M. Dorget, J.-F. Palierne, and A. Pouchelon, Shear elasticity and yield stress of silica–silicone physical gels: Fractal approach, Journal of Rheology 43, 305 (1999).
  • Sztucki and Narayanan [2007] M. Sztucki and T. Narayanan, Development of an ultra-small-angle X-ray scattering instrument for probing the microstructure and the dynamics of soft matter, Journal of Applied Crystallography 40, s459 (2007).
  • Livsey [1987] I. Livsey, Neutron scattering from concentric cylinders. intraparticle interference function and radius of gyration, Journal of the Chemical Society, Faraday Transactions 2: Molecular and Chemical Physics 83, 1445 (1987).
  • Murata et al. [1994] K. Murata, M. Aoki, T. Suzuki, T. Harada, H. Kawabata, T. Komori, F. Ohseto, K. Ueda, and S. Shinkai, Thermal and light control of the sol-gel phase transition in cholesterol-based organic gels. novel helical aggregation modes as detected by circular dichroism and electron microscopic observation, Journal of the American Chemical Society 116, 6664 (1994).
  • Sarkar et al. [2020] A. Sarkar, R. Sasmal, C. Empereur-Mot, D. Bochicchio, S. V. Kompella, K. Sharma, S. Dhiman, B. Sundaram, S. S. Agasti, G. M. Pavan, et al., Self-sorted, random, and block supramolecular copolymers via sequence controlled, multicomponent self-assembly, Journal of the American Chemical Society 142, 7606 (2020).
  • Lai et al. [2016] Z. Lai, Y. Chen, C. Tan, X. Zhang, and H. Zhang, Self-assembly of two-dimensional nanosheets into one-dimensional nanostructures, Chem 1, 59 (2016).
  • Landman et al. [2018] J. Landman, S. Ouhajji, S. Prévost, T. Narayanan, J. Groenewold, A. P. Philipse, W. K. Kegel, and A. V. Petukhov, Inward growth by nucleation: Multiscale self-assembly of ordered membranes, Science advances 4, eaat1817 (2018).
  • Monk et al. [1999] P. M. Monk, N. M. Hodgkinson, and R. D. Partridge, The colours of charge-transfer complexes of methyl viologen: effects of donor, ionic strength and solvent, Dyes and pigments 43, 241 (1999).
  • Yoshikawa and Nishikiori [2005] H. Yoshikawa and S.-i. Nishikiori, Crystal structures and spectroscopic properties of polycyano–polycadmate host clathrates including a ct complex guest of methylviologen dication and aromatic donor, Dalton Transactions , 3056 (2005).
@@
  • De Greef et al. [2009] T. F. De Greef, M. M. Smulders, M. Wolffs, A. P. Schenning, R. P. Sijbesma, and E. W. Meijer, Supramolecular polymerization, Chemical Reviews 109, 5687 (2009).
  • Mao et al. [2016] B. Mao, T. Divoux, and P. Snabre, Normal force controlled rheology applied to agar gelation, Journal of Rheology 60, 473 (2016).
  • Yuan et al. [2011] B. Yuan, X.-Y. Liu, J.-L. Li, and H.-Y. Xu, Volume confinement induced microstructural transitions and property enhancements of supramolecular soft materials, Soft Matter 7, 1708 (2011).
@@
  • Liu et al. [2014] Y. Liu, R. Y. Wang, J. L. Li, B. Yuan, M. Han, P. Wang, and X. Y. Liu, Identify kinetic features of fibers growing, branching, and bundling in microstructure engineering of crystalline fiber network, CrystEngComm 16, 5402 (2014).

### v2

@@ IV Conclusion
In conclusion, this study sheds light on the formation and characteristics of viologen-based supramolecular crystal gels formed by self-assembly of low molecular weight gelators into interconnected crystalline fibers, resulting in a three-dimensional soft solid network. Our in-depth analyzes of the gelation kinetics triggered by thermal quenching reveal the mechanism at play to form the 3D network. We first confirm that the viologen-based gelator $Chol_{2}V^{2+}$ crystallizes into nanometer-radius hollow tubes, which further assemble into micro to millimetric spherulites. Then, we demonstrate the spherulite growth is consistent with the Avrami theory. Local characterization of the growth kinetics using optical microscopy and USALS techniques allowed us to determine independently the exponents in the Avrami model (Eq. 2: $n=n_{N}+dn_{G}$). It reveals that the crystallization process is driven by pre-existing nuclei ($n_{N}=0$), with growth controlled by interfaces ($n_{G}=1$) and characterized by branching resulting in the formation of fractal spherulites of fractal dimension $d$. Finally, our results demonstrate the possibility of adjusting the gel properties by varying the quenching temperature. Lowering the quenching temperature leads to the formation of denser and smaller spherulites. We establish that the elastic properties of the gel cannot be described by the model proposed in [27, 28, 30]. We speculate that the elasticity of the gel stems from both the spherulite elasticity and the way they connect to one another: the spherulites densification obtained at low quenching temperature occurs at the expense of the connectivity between spherulite. This comprehensive analysis provides valuable insights into the design and manipulation of supramolecular crystal gels.
@@ V.2 Radius of the spherulites over time
In order to characterize the formation of ${Chol_{2}V^{2+}/I^{-}}$ CT complexes in the absence of any other anionic species the viologen salt $Chol_{2}VI_{2}$ was prepared using the following protocol. Compound $Chol_{2}VBr_{2}$ (0.410 g, 0.32 mmol), $LiTFSI$ (3.494 g, 6.4 mmol) and 400 mL technical $EtOH$ were introduced in a 1L round bottom flask and refluxed overnight under stirring.  80% of the solvent was evaporated under reduced pressure, 200mL water were added and the mix was left in the fridge overnight. The mix was filtered on and washed with 3x40mL water. The white solid was then dried under vacuum, 420mg of colorless solid was recovered. This solid was dissolved in 20 mL acetone, Tetrabuthylammonium iodide (820 mg, 2.2mmol) were added and the mix was stirred at room temperature for 1h. The mix was filtered and washed with acetone (3x20mL). $Chol_{2}VI_{2}$ was obtained as an orange solid in 73% yield. The evolution of the UV-Vis absorption spectrum of $Chol_{2}VI_{2}$ is displayed in Fig. 9(a). In Fig. 9(b), we observe that the absorption at $\lambda=440$ nm increase linearly with $Chol_{2}VI_{2}$ concentration leading to an adsorption coefficient is $\epsilon_{\lambda=450nm}^{path=1cm}=6.3\pm 0.3\times 10^{3}$ mol/L/cm.
Figure 9: a) Absorption spectra obtained at 20°C from $Chol_{2}VI_{2}$ solutions and gels in pentanol (2mm quartz cuvette). Samples were prepared after applying a heating cooling cycle to $Chol_{2}VI_{2}$ suspensions in 1-Pentanol. b) Evolution of the Absorbance at 440 nm with the concentration in $Chol_{2}VI_{2}$. The continuous red line to the best linear fit.
@@ Acknowledgment
  • [1] A. Ajayaghosh, V. K. Praveen, and C. Vijayakumar. Organogels as scaffolds for excitation energy transfer and light harvesting. Chemical Society Reviews, 37(1):109–122, 2008.
  • [2] V. Andrieux, T. Gibaud, J. Bauland, T. Divoux, S. Manneville, S. Guy, A. Bensalah-Ledoux, L. Guy, F. Chevallier, D. Frath, and C. Bucher. Chiral and conductive viologen-based supramolecular gels exhibiting tunable charge-transfer properties. J. Mater. Chem. C, 11:12764–12775, 2023.
@@
  • [3] M. Avrami. Granulation, phase change, and microstructure kinetics of phase change. iii. The Journal of chemical physics, 9(2):177–184, 1941.
  • [4] E. Carretti, M. Bonini, L. Dei, B. H. Berrie, L. V. Angelova, P. Baglioni, and R. G. Weiss. New frontiers in materials science for art conservation: responsive gels and beyond. Accounts of chemical research, 43(6):751–760, 2010.
  • [5] H. Chen, V. Brasiliense, J. Mo, L. Zhang, Y. Jiao, Z. Chen, L. O. Jones, G. He, Q.-H. Guo, X.-Y. Chen, et al. Single-molecule charge transport through positively charged electrostatic anchors. Journal of the American Chemical Society, 143(7):2886–2895, 2021.
  • [6] S. Chowdhury, Q. Reynard-Feytis, C. Roizard, D. Frath, F. Chevallier, C. Bucher, and T. Gibaud. Light-Controlled Aggregation and Gelation of Viologen-Based Coordination Polymers. Journal of Physical Chemistry B, 125(43):12063–12071, 2021.
  • [7] S. Datta and S. Bhattacharya. Carbon-nanotube-mediated electrochemical transition in a redox-active supramolecular hydrogel derived from viologen and an l-alanine-based amphiphile. Chemistry – A European Journal, 22(22):7524–7532, 2016.
  • [8] S. Datta, N. Dey, and S. Bhattacharya. Electrochemical probing of hydrogelation induced by the self-assembly of a donor–acceptor complex comprising pyranine and viologen. Chem. Commun., 53:2371–2374, 2017.
@@
  • [9] T. F. De Greef, M. M. Smulders, M. Wolffs, A. P. Schenning, R. P. Sijbesma, and E. W. Meijer. Supramolecular polymerization. Chemical Reviews, 109(11):5687–5754, 2009.
  • [10] S. Dhiman, K. Jalani, and S. J. George. Redox-mediated, transient supramolecular charge-transfer gel and ink. ACS Applied Materials & Interfaces, 12(5):5259–5264, 2020.
  • [11] G. I. Dzhardimalieva, B. C. Yadav, S. Singh, and I. E. Uflyand. Self-healing and shape memory metallopolymers: state-of-the-art and future perspectives. Dalton Trans., 49:3042–3087, 2020.
  • [12] W. Fang, Y. Zhang, J. Wu, C. Liu, H. Zhu, and T. Tu. Recent advances in supramolecular gels and catalysis. Chemistry – An Asian Journal, 13(7):712–729, 2018.
  • [13] C. Gila-Vilchez, M. C. Mañas-Torres, J. A. González-Vera, F. Franco-Montalban, J. A. Tamayo, F. Conejero-Lara, J. M. Cuerva, M. T. Lopez-Lopez, A. Orte, and L. Á. de Cienfuegos. Insights into the co-assemblies formed by different aromatic short-peptide amphiphiles. Polymer Chemistry, 12(47):6832–6845, 2021.
  • [14] Y. Han, C. Nickle, Z. Zhang, H. P. Astier, T. J. Duffin, D. Qi, Z. Wang, E. Del Barco, D. Thompson, and C. A. Nijhuis. Electric-field-driven dual-functional molecular switches in tunnel junctions. Nature materials, 19(8):843–848, 2020.
  • [15] M. C. Heijna, M. J. Theelen, W. J. van Enckevort, and E. Vlieg. Spherulitic growth of hen egg-white lysozyme crystals. The Journal of Physical Chemistry B, 111(7):1567–1573, 2007.
  • [16] C. D. Jones and J. W. Steed. Gels with sense: Supramolecular materials that respond to heat, light and sound, 2016.
  • [17] C. Kahlfuss, T. Gibaud, S. Denis-Quanquin, S. Chowdhury, G. Royal, F. Chevallier, E. Saint-Aman, and C. Bucher. Redox-Induced Molecular Metamorphism Promoting a Sol/Gel Phase Transition in a Viologen-Based Coordination Polymer. Chemistry - A European Journal, 24(49):13009–13019, 2018.
  • [18] K. K. Kartha, S. S. Babu, S. Srinivasan, and A. Ajayaghosh. Attogram sensing of trinitrotoluene with a self-assembled molecular gelator. Journal of the american chemical society, 134(10):4834–4841, 2012.
  • [19] M. Kathiresan, B. Ambrose, N. Angulakshmi, D. E. Mathew, D. Sujatha, and A. M. Stephan. Viologens: a versatile organic molecule for energy storage applications. Journal of Materials Chemistry A, 9(48):27215–27233, 2021.
  • [20] Z. Lai, Y. Chen, C. Tan, X. Zhang, and H. Zhang. Self-assembly of two-dimensional nanosheets into one-dimensional nanostructures. Chem, 1(1):59–77, 2016.
  • [21] V. Lakshminarayanan, C. Chockalingam, E. Mendes, and J. H. van Esch. Gelation kinetics-structure analysis of ph-triggered low molecular weight hydrogelators. ChemPhysChem, 22(21):2256–2261, 2021.
  • [22] R. Lam, L. Quaroni, T. Pedersen, and M. A. Rogers. A molecular insight into the nature of crystallographic mismatches in self-assembled fibrillar networks under non-isothermal crystallization conditions. Soft Matter, 6(2):404–408, 2010.
  • [23] J. Landman, S. Ouhajji, S. Prévost, T. Narayanan, J. Groenewold, A. P. Philipse, W. K. Kegel, and A. V. Petukhov. Inward growth by nucleation: Multiscale self-assembly of ordered membranes. Science advances, 4(6):eaat1817, 2018.
  • [24] M. Lescanne, A. Colin, O. Mondain-Monval, F. Fages, and J.-L. Pozzo. Structural aspects of the gelation process observed with low molecular mass organogelators. Langmuir, 19(6):2013–2020, 2003.
  • [25] J. Li, S. Pudar, H. Yu, S. Li, J. S. Moore, J. Rodríguez-López, N. E. Jackson, and C. M. Schroeder. Reversible switching of molecular conductance in viologens is controlled by the electrochemical environment. The Journal of Physical Chemistry C, 125(40):21862–21872, 2021.
  • [26] J. L. Li and X. Y. Liu. Architecture of supramolecular soft functional materials: From understanding to micro-/nanoscale engineering. Advanced Functional Materials, 20(19):3196–3216, 2010.
  • [27] X. Y. Liu and P. D. Sawant. Formation kinetics of fractal nanofiber networks in organogels. Applied Physics Letters, 79(21):3518–3520, 2001.
  • [28] X. Y. Liu and P. D. Sawant. Determination of the fractal characteristic of nanofiber-network formation in supramolecular materials. ChemPhysChem, 3(4):374–377, 2002.
  • [29] X. Y. Liu, P. D. Sawant, W. B. Tan, I. Noor, C. Pramesti, and B. Chen. Creating new supramolecular materials by architecture of three-dimensional nanocrystal fiber networks. Journal of the American Chemical Society, 124(50):15055–15063, 2002.
@@
  • [30] Y. Liu, R. Y. Wang, J. L. Li, B. Yuan, M. Han, P. Wang, and X. Y. Liu. Identify kinetic features of fibers growing, branching, and bundling in microstructure engineering of crystalline fiber network. CrystEngComm, 16(24):5402–5408, 2014.
  • [31] I. Livsey. Neutron scattering from concentric cylinders. intraparticle interference function and radius of gyration. Journal of the Chemical Society, Faraday Transactions 2: Molecular and Chemical Physics, 83(8):1445–1452, 1987.
  • [32] Z. Ma, G. Zhang, X. Zhai, L. Jin, X. Tang, M. Yang, P. Zheng, and W. Wang. Fractal crystal growth of poly (ethylene oxide) crystals from its amorphous monolayers. Polymer, 49(6):1629–1634, 2008.
  • [33] B. Mao, T. Divoux, and P. Snabre. Normal force controlled rheology applied to agar gelation. Journal of Rheology, 60(3):473–489, 2016.
  • [34] V. H. Mareau and R. E. Prud’Homme. In-situ hot stage atomic force microscopy study of poly ($\varepsilon$-caprolactone) crystal growth in ultrathin films. Macromolecules, 38(2):398–408, 2005.
  • [35] P. M. Monk, N. M. Hodgkinson, and R. D. Partridge. The colours of charge-transfer complexes of methyl viologen: effects of donor, ionic strength and solvent. Dyes and pigments, 43(3):241–251, 1999.
  • [36] K. Murata, M. Aoki, T. Suzuki, T. Harada, H. Kawabata, T. Komori, F. Ohseto, K. Ueda, and S. Shinkai. Thermal and light control of the sol-gel phase transition in cholesterol-based organic gels. novel helical aggregation modes as detected by circular dichroism and electron microscopic observation. Journal of the American Chemical Society, 116(15):6664–6676, 1994.
  • [37] P. Nasr, H. Leung, F.-I. Auzanneau, and M. A. Rogers. Supramolecular fractal growth of self-assembled fibrillar networks. Gels, 7(2):46, 2021.
  • [38] G. M. Newbloom, K. M. Weigandt, and D. C. Pozzo. Electrical, mechanical, and structural characterization of self-assembly in poly(3-hexylthiophene) organogel networks. Macromolecules, 45(8):3452–3462, 2012.
  • [39] Q. V. Nguyen, P. Martin, D. Frath, M. L. Della Rocca, F. Lafolet, S. Bellinck, P. Lafarge, and J.-C. Lacroix. Highly efficient long-range electron transport in a viologen-based molecular junction. Journal of the American Chemical Society, 140(32):10131–10134, 2018.
  • [40] M. Pernetti, K. van Malssen, D. Kalnin, and E. Flöter. Structuring edible oil with lecithin and sorbitan tri-stearate. Food Hydrocolloids, 21(5-6):855–861, 2007.
  • [41] J.-M. Piau, M. Dorget, J.-F. Palierne, and A. Pouchelon. Shear elasticity and yield stress of silica–silicone physical gels: Fractal approach. Journal of Rheology, 43(2):305–314, mar 1999.
  • [42] K. V. Rao, K. Jayaramulu, T. K. Maji, and S. J. George. Supramolecular hydrogels and high-aspect-ratio nanofibers through charge-transfer-induced alternate coassembly. Angewandte Chemie International Edition, 49(25):4218–4222.
  • [43] C. Roizard, V. Andrieux, S. A. Shehimy, S. Chowdhury, Q. Reynard-Feytis, C. Kahlfuss, E. Saint-Aman, F. Chevallier, C. Bucher, T. Gibaud, and D. Frath. Photoredox processes in the aggregation and gelation of electron-responsive supramolecular polymers based on viologen. ECS Advances, 1(2):020502, 2022.
  • [44] A. Sarkar, R. Sasmal, C. Empereur-Mot, D. Bochicchio, S. V. Kompella, K. Sharma, S. Dhiman, B. Sundaram, S. S. Agasti, G. M. Pavan, et al. Self-sorted, random, and block supramolecular copolymers via sequence controlled, multicomponent self-assembly. Journal of the American Chemical Society, 142(16):7606–7617, 2020.
  • [45] L. Saunders and P. X. Ma. Self-healing supramolecular hydrogels for tissue engineering applications. Macromolecular bioscience, 19(1):1800313, 2019.
  • [46] J. H. Shi, X. Y. Liu, J. L. Li, C. S. Strom, and H. Y. Xu. Spherulitic networks: from structure to rheological property. The journal of physical chemistry b, 113(14):4549–4554, 2009.
  • [47] K. Shirzad and C. Viney. A critical review on applications of the avrami equation beyond materials science. Journal of the Royal Society Interface, 20(203):20230242, 2023.
  • [48] A. G. Shtukenberg, Y. O. Punin, E. Gunn, and B. Kahr. Spherulites. Chemical reviews, 112(3):1805–1838, 2012.